1. Fifteen years ago, you put away $7,500. Today, that investment is now worth $22,068. What is the average annual rate of return you earned on your investment?

2. Your company takes out a 10-year term loan to borrow $1,325,000 to finance a building expansion. The annual interest rate on the loan is 4.75% and equal payments are made monthly. How much will the monthly payment be?                     

3. Referring back to question #2, how much total interest will the company be paying over the life of the term loan?
Answer #1

Ans 1) 7.46%

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r%)^n
22068 / 7500 = (1 + r%)^15
2.9424 = (1 + r%)^15
r = 7.46%